Anthony is 5 years older than his sister Ellen, and 7 years ago the sum of their ages was 23 years. In how many years will Anthony be 30 years old?

Answer by ikleyn(52785)

.
Anthony is 5 years older than his sister Ellen, and 7 years ago the sum of their ages was 23 years.
In how many years will Anthony be 30 years old?
~~~~~~~~~~~~~~~~~~~~~~~~~~

The condition says

A - E = 5,            (1)    ("Anthony is 5 years older than his sister Ellen")
(A-7) + (E-7) = 23.   (2)    ("7 years ago the sum of their ages was 23 years. ")

Simplify and rewrite equivalently this system in the form:

A - E =  5,           (3)
A + E = 37.           (4)

Now add the equations (3) and (4). You will get

2A = 42.

Can you complete the solution from this point on your own?

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".